# Anisotropic Spin Ice on a Breathing Pyrochlore Lattice

### v1

@@ III.6 Ferromagnetic, $\delta_{A},\delta_{B}<0$
In the case $\delta_{A},\delta_{B}<0$, the total charge is $\mathsf{Q}_{A/B}=0$ and the partial charges are $\mathsf{q}_{A/B}^{03}=-\mathsf{q}_{A/B}^{12}=\pm 2$. On each tetrahedron, there are two possible ground states – the ones with magnetization vector parallel or antiparallel with the global $z$ axis. By fixing a single spin on a single tetrahedron, however, we will fix all the spins in the lattice, as the tetrahedra are corner-sharing. So, in this case, the total ground state degeneracy is only $2$. The spin configuration on all tetrahedra of the same kind is equal. However, the magnetization vector of tetrahedra of different kinds will still point in the same direction, either into $+\hat{z}$ or $-\hat{z}$. This means that this limiting case can be seen as a ferromagnet, where the ground state spontaneously breaks the symmetry of the Hamiltonian, as pointed out in [17]. The ground state of the system is the fully polarized state, where the local magnetization vector of all tetrahedra points in the same direction. For this reason, we refer to this phase as ferromagnetic.
In summary, here we have a system with a double degenerate ferromagnetic ground state. We can change the ground state by flipping all the spins of the system simultaneously. Here, the $\mathbb{Z}_{2}$ symmetry of the Hamiltonian is broken spontaneously.
@@ IV Thermodynamic properties
For any finite value of $\delta_{A/B}$, the full model’s ground state degeneracy is reduced compared to the isotropic spin ice case. We can see that, depending on the respective sign of the two distortions, the ground state degeneracy can either scale exponentially with the full system volume (as in the planar ice case) – or with just a lower-dimensional cut of the system. When considering how to generate all possible ground state configurations within a single phase, we see that in all cases, lower dimensional cuts of the system decouple and can be changed individually.
@@
where $S_{\infty}=N\log(2)$ is the infinite temperature entropy stemming from the $2^{N}$ possible spin orientations. These measures are indicative of the system’s equilibration properties. Numerical data is obtained via Monte Carlo simulations, where updates are performed through attempted single spin-flips as well as worm loop updates and cluster updates to accelerate equilibration. Worm loop updates consist of finding a closed loop of spins aligned head to tail and flipping all spins part of the loop [41, 42, 43]. The closed loop cannot create or destroy monopoles but can create and move partial charges. The flip is then accepted with the Metropolis probability of $\text{min}[1,\exp{(-\Delta E/T)}]$. Since, for the anisotropic systems, the acceptance probability of a random worm loop is drastically reduced when entering into the restricted ground state manifold, additional cluster updates are employed. These cluster updates are specifically tailored to the ground state-generating operator structure. A detailed description is given in App. E. The system is first equilibrated at a specific temperature before measurements are taken. A measurement series at a specific temperature is completed when $N_{\text{sweeps}}=N_{\text{spin}}\cdot 10^{3}$ update-attempts are accepted. All further details of the numerical data are given in App. D. The data presented in Fig. 3 is obtained for $L_{x}=L_{y}=L_{z}=4$, corresponding to $N_{\text{spin}}=1024$ or $L_{x}\times L_{y}\times L_{z}=64$ face-centered cubic (fcc) unit cells.
We will discuss specific heat and entropy for the isotropic spin ice case, $\delta_{A/B}=0$, as well as well as five specific points in the phase diagram. We will touch on the generalization to all cases at the end of the section. To compare the characteristics of the five different cases, we fix the values of $\delta_{A},\delta_{B}=\delta\cdot\{+1,0,-1\}$. We will also assume that the exchange interaction on both sublattices is equally set to $J_{A}=J_{B}=1$, and $\delta_{i}<J_{i}$, such that the exchange interaction is still the dominant interaction. The latter means the order of excitations on top of the ground state is fixed such that a monopole excitation ($\mathsf{Q}=\pm 2$) is energetically more costly than a partial-charge change ($\Delta\mathsf{q}^{\alpha\beta}=\pm 2$). The data presented in Fig. 3 is for $\delta=0.05$.
Figure 3: Specific heat and entropy density of anisotropic spin ice. (a) While the isotropic case shows only one bump in the specific heat, all anisotropic cases show a second bump at a lower temperature. (b) Zoom in on the second bump in the specific heat. (c) Entropy density for all anisotropic cases. The entropy density of isotropic spin ice approaches the Pauling residual entropy value, while for all other anisotropic cases, the entropy density is lowered. Residual entropy density, $s_{\text{res}}=S(T=0)/N$, is shown in the inset. For isotropic spin ice and the planar ice case, the residual entropy stays finite, while for the other anisotropic cases, the residual entropy drops with increasing system size. All data is obtained for $L_{x}=L_{y}=L_{z}=4$ face-centered cubic (fcc) unit cells, corresponding to $N_{\text{spin}}=1024$.
@@ IV.1 Isotropic case
At high temperatures, spin ice is in a trivially disordered state as thermal fluctuations dominate. Here, the system is in a paramagnetic regime, and no correlations in the spin orientations can be found. Upon cooling, however, the system enters into the spin ice manifold. Here, each tetrahedron tends towards satisfying the “two in - two out” rule to locally minimize the energy. This crossover can be found, according to extensive numerical and experimental studies [44, 45, 3, 46, 47, 1] at around $T_{c}\approx J$. Above this crossover, thermal fluctuations easily introduce new monopoles. This crossover can be experimentally observed by a Schottky-like peak in the specific heat. With simple considerations for a single tetrahedron, neglecting double monopoles, we can confirm this estimate for the point of this crossover: there are six possibilities for a ground state on a tetrahedron, while there are eight configurations with a single monopole. Calculating the maximum of the specific heat obtained by the derivative of $Z=6+8\cdot e^{-2\beta J}$ gives $T_{c}\approx 0.802J$. Following eq. (4) also the entropy at a specific temperature can be computed. It can be observed that for temperatures below this crossover regime, the entropy density approaches a finite value of $s_{\text{res}}=S/(k_{B}N)\approx\frac{1}{2}\ln\left(\frac{3}{2}\right)$, the Pauling residual entropy [37]. This has been reported in experiments as well, for example in Dy2Ti2O7 [3], indicating the large degeneracy of ground states. In Fig. 3, the isotropic spin ice is the grey curve that follows the expected behavior well.
@@ IV.2 Anisotropic Cases
The specific heat and entropy in the temperature regime from $T=0.08J-15.0J$ for the five different anisotropic phases are shown in Fig. 3. For all of our analyzed cases, we can observe that the specific heat follows the behavior of spin ice until the crossover at $T\approx J$. This can be understood as all the systems having similar behavior at high temperatures: a paramagnetic regime where thermal fluctuations govern the system. At around the first crossover, the behavior of the isotropic and the anisotropic systems still match. Here, the system crosses over from a paramagnetic regime to a regime, where spins on a tetrahedron follow the ice rule. However, below this point, the behavior of the different phases is very different. In fact, we can observe either a second crossover (in the planar ice phase), indicated by a second bump in the specific heat, or a phase transition toward an ordered phase, indicated by a divergence in the specific heat. This allows us to make a distinction between the symmetry-broken phases and the non-symmetry-broken phases. For all systems, we can observe that the entropy density is lower than the Pauling estimate. Next we analyze the five different phases in more detail and distinguish between the non-symmetry-broken phases and the symmetry-broken phases at finite temperatures.
@@ IV.2.1 Non-symmetry-broken phases
If, the $\delta$-parameters on the two sublattices were different, $\delta_{A}\neq\delta_{B}>0$, then we will observe a first peak around $T_{\delta_{A}}\approx 2\delta_{A}$ and a second peak around $T_{\delta_{B}}\approx 2\delta_{B}$. If the two are equal, the peaks approximately stack in their common crossover regime. We can again see the entropy density drop below the Pauling estimate, approaching zero for increasing system sizes as $s_{\text{res}}=\frac{1}{4\cdot L_{x}L_{y}}\ln{(2^{\text{GCF}(L_{x},L_{y})})}$ ).
@@ IV.2.2 Symmetry-broken phases
If on any sublattice the strain-parameter $\delta_{A}/\delta_{B}<0$, a phase transition into a symmetry-broken state will occur at a finite temperature. The phase transition presents itself as a divergence in the specific heat. Due to finite size effects, the divergence will appear as a narrow peak with a finite height in Fig. 3.
In the omni-plane case, $\delta_{A}<0,\delta_{B}=0$, the spins on the $A$ tetrahedra will order in such a manner that the local magnetization vector of the tetrahedron is $M\parallel\hat{z}$. We can now distinguish between three types of ground states: the $A$-tetrahedra will either order antiferromagnetically in all individual planes perpendicular to the $z$ direction, or they will order ferromagnetically in all individual planes perpendicular to the $x$ or $y$ direction. In all cases, the $\mathbb{Z}_{2}$ symmetry in the respective plane is broken, and therefore, a phase transition will appear. The critical temperature is approximately $T_{c}\approx\frac{4\delta}{\log{(6)}}$. Due to the phase transition, the entropy drops sharply below the critical temperature.
In the plane-paramegntic case, $\delta_{A}>0$ and $\delta_{B}<0$, the local magnetization vector of all $B$ tetrahedra are $\mathbf{M}_{i}\parallel\hat{z}$. Within single $x,y$-planes of $B$ tetrahedra, the $B$ tetrahedra form a square lattice. Due to the order of the spin-pairs on $A$-tetrahedra, the $B$ tetrahedra will order antiferromagnetically within the planes and break the planes $\mathbb{Z}_{2}$ symmetry. This transition to antiferromagnetic order in the decoupled planes will occur at a critical temperature of around $T_{c}\approx\frac{4\delta}{\log{(3)}}$. At the critical point, the specific heat is expected to diverge, and a steep drop in entropy is visible.
For the ferromagnetic case, where $\delta_{A}=\delta_{B}<0$, it has been shown that a transition akin to the one found in hydrogen-bonded ferroelectrics known as the KDP (potassium dihydrogen phosphate) model [49, 50] takes place, ordering the system ferromagnetically. Here, a sharp divergence of the specific heat, as well as a sharp dip in the entropy is visible [17, 18]. We can calculate the exact critical temperature by considering the energy cost of a string excitation and the entropy of this excitation [28]. Let us consider a ground state of the system. All the spins are equally aligned on all tetrahedra, such that the global magnetization vector of all tetrahedra points in the same direction. A string excitation, now, consists of flipping a row of spins in the $z$ direction and thereby increasing the energy. Since we are introducing a partial charge in 4 tetrahedra per unit cell, for $L_{z}$ unit cells in $z$ direction, the energy for such an excitation is $\Delta U=4\delta\cdot 4L_{z}=16\delta L_{z}$. Because there are two possible ways for a string to enter each tetrahedron, and we flip $4L_{z}$ tetrahedra, the entropic gain of a string excitation is $\Delta S=\ln(2^{4L_{z}})$. So the difference in free energy for a string excitation is
@@
In summary, in all cases, the specific heat has a second signature below the first cross-over at $T\approx J$. If on any sublattice there is a negative strain parameter, $\delta_{i}<0$, we expect a phase transition at a finite temperature into a symmetry-broken phase. The symmetry breaking will, however, occur only in dimensionally reduced subsystems if one of the strain parameters is non-negative. Here, the order parameter is defined on a dimensionally reduced support – a two-dimensional plane – and the symmetry that is broken is the intermediate symmetry of the plane [32, 33, 31]. This also leads to a vanishing entropy density. If both strain parameters are positive, one-dimensional lines will effectively decouple, leading to antiferromagnetic order on lines at zero temperature. Similar behavior has been observed in the anisotropic antiferromagnetic triangular lattice Ising model, where strong coupling along one lattice direction leads to effective dimensional reduction and frustrated inter-chain interactions. Despite the two-dimensional connectivity, the system may exhibit behavior akin to weakly coupled one-dimensional chains without a finite-temperature phase transition [51]. If only one strain parameter is non-zero and positive, a cross-over into a regime where the two-dimensional planes decouple with an extensive ground state degeneracy occurs. Within planes, the planar ice constraint holds. This is the only phase where the entropy density stays finite even for infinite system size.
@@
Here, we assume the systems to be at equilibrium. To reach thermal equilibrium in numerical simulations, techniques such as cluster updates or worm loop updates are employed to overcome large energy barriers. In fact, we can observe that the acceptance rate for single spin-flips is reduced significantly below the first crossover into the spin ice regime at $T\approx J$. For isotropic spin ice, the worm algorithm has been proven useful to overcome the long waiting times to overcome the energy barriers and equilibrate spin ice systems [15, 41, 42, 43]. The acceptance rate for worm loop updates is exactly one below the crossover into the spin ice manifold for the isotropic case. However, for the anisotropic systems, even the acceptance rate for the worm algorithm reduces close to the crossover into the respective restricted manifold. This renders equilibrating the anisotropic systems numerically very hard, as updates are correlated with extremely long waiting times that increase with increasing system size. It is, however, important to point out that these are higher-order processes that will also take extremely long in an experimental setup. It is expected that cooling the system too quickly will hide these features, and the system might then seem to behave akin to spin ice [9].
@@ V Spin Structure Factor
In a scattering event with neutron polarization parallel to $\hat{z}_{\text{sc}}$, one can define the orthonormal basis vectors for each scattering wave vector $\mathbf{q}\perp\hat{z}_{\text{sc}}$ as $\hat{x}_{\text{sc}}\equiv\hat{\mathbf{q}}$ and $\hat{y}_{\text{sc}}\equiv\hat{z}_{\text{sc}}\times\hat{x}_{\text{sc}}$. The spin vector is defined as $\hat{S}^{z}_{i,\alpha}=S^{z}_{i,\alpha}\cdot\hat{e}_{i}$, where $\alpha$ is the index referring to the four sites in the face-centered cubic (fcc) unit cell, while $i$ is the spin index, referring to the site on a single tetrahedron. The unit vectors $\hat{e}_{i}$ are the local basis vectors pointing into the center of a single tetrahedron, as defined in App. B.
@@
In polarized neutron scattering experiments, with $\hat{z}_{\text{sc}}\parallel[1\bar{1}0]$, the spin ice material Ho2Ti2O7 and Dy2Ti2O7 show particular pinch points in the $[hhl]$ plane in the SF channel, indicative of the dipolar form of the spin correlations in three dimensions, while in the NSF channel no particular features can be seen [2, 52]. Now, we want to understand how this structure factor will differ in the anisotropic phases compared to isotropic spin ice. Numerically, the spin structure factor can be obtained by sampling states via a Monte Carlo procedure or by an analytical mean-field analysis. We will employ both techniques and show the correspondence of the data in Fig. 4.

### v2

@@ III.6 Ferromagnetic, $\delta_{A},\delta_{B}<0$
In the case $\delta_{A},\delta_{B}<0$, the total charge is $\mathsf{Q}_{A/B}=0$ and the partial charges are $\mathsf{q}_{A/B}^{03}=-\mathsf{q}_{A/B}^{12}=\pm 2$. On each tetrahedron, there are two possible ground states – the ones with magnetization vector parallel or antiparallel to the global $z$ axis. By fixing a single spin on a single tetrahedron, however, we will fix all the spins in the lattice, as the tetrahedra are corner-sharing. So, in this case, the total ground state degeneracy is only $2$. The spin configuration on all tetrahedra of the same kind is equal. However, the magnetization vector of tetrahedra of different kinds will still point in the same direction, either into $\hat{z}$ or $\hat{z}$. This means that this limiting case can be seen as a ferromagnet, where the ground state spontaneously breaks the symmetry of the Hamiltonian, as pointed out in [17]. The ground state of the system is the fully polarized state, where the local magnetization vector of all tetrahedra points in the same direction. For this reason, we refer to this phase as ferromagnetic. Also in this case, partial charge excitations and monopole excitations are immobile, since crossing any tetrahedron necessarily increases the energy of the system by $4|\delta|$.
In summary, here we have a system with a double-degenerate ferromagnetic ground state. We can change the ground state by flipping all the spins of the system simultaneously. Here, the $\mathbb{Z}_{2}$ symmetry of the Hamiltonian is broken spontaneously.
@@ IV Thermodynamic properties
For any finite value of $\delta_{A/B}$, the full model’s ground state degeneracy is reduced compared to the isotropic spin ice case. We can see that, depending on the respective sign of the two distortions, the ground state degeneracy can either scale exponentially with the full system volume (as in the planar ice case) – or with just a lower-dimensional cut of the system. When considering how to generate all possible ground state configurations within a single phase, we see that in all cases, lower-dimensional cuts of the system decouple and can be changed individually.
@@
$S(T)=S_{\infty}-\int_{T}^{\infty}dT^{\prime}\frac{C_{V}(T^{\prime})}{T^{\prime}},$ (4)
@@
We will discuss specific heat and entropy for the isotropic spin ice case, $\delta_{A/B}=0$, as well as five specific points in the phase diagram. We will touch on the generalization to all cases at the end of the section. To compare the characteristics of the five different cases, we fix the values of $\delta_{A},\delta_{B}=\delta\cdot\{+1,0,-1\}$. We will also assume that the exchange interaction on both sublattices is equally set to $J_{A}=J_{B}=1$, and $\delta_{i}<J_{i}$, such that the exchange interaction is still the dominant interaction. The latter means the order of excitations on top of the ground state is fixed such that a monopole excitation ($\mathsf{Q}=\pm 2$) is energetically more costly than a partial-charge change ($\Delta\mathsf{q}^{\alpha\beta}=\pm 2$). The data presented in Fig. 3 is for $\delta=0.05$.
[…]
@@ IV.1 Isotropic case
At high temperatures, spin ice is in a trivially disordered state as thermal fluctuations dominate. Here, the system is in a paramagnetic regime, and no correlations in the spin orientations can be found. Upon cooling, however, the system enters into the spin ice manifold. Here, each tetrahedron tends towards satisfying the “two in - two out” rule to locally minimize the energy. This crossover can be found, according to extensive numerical and experimental studies [58, 59, 3, 38, 60, 1] at around $T_{c}\approx J$. Above this crossover, thermal fluctuations easily introduce new monopoles. This crossover can be experimentally observed by a Schottky-like peak in the specific heat. With simple considerations for a single tetrahedron, neglecting double monopoles, we can confirm this estimate for the point of this crossover: there are six possibilities for a ground state on a tetrahedron, while there are eight configurations with a single monopole. Calculating the maximum of the specific heat obtained by the derivative of $Z=6+8\cdot e^{-2\beta J}$ gives $T_{c}\approx 0.802J$. Following eq. (4), also, the entropy at a specific temperature can be computed. It can be observed that for temperatures below this crossover regime, the entropy density approaches a finite value of $s_{\text{res}}=S/(k_{B}N)\approx\frac{1}{2}\ln\left(\frac{3}{2}\right)$, the Pauling residual entropy [49]. This has been reported in experiments as well, for example, in Dy2Ti2O7 [3], indicating the large degeneracy of ground states. In Fig. 3, the isotropic spin ice is the grey curve that follows the expected behavior well.
@@ IV.2 Anisotropic Cases
The specific heat and entropy in the temperature regime from T=0.08J15.0JT=0.08J-15.0J for the five different anisotropic phases are shown in Fig. 3. For all of our analyzed cases, we can observe that the specific heat follows the behavior of spin ice until the crossover at $T\approx J$. This can be understood as all the systems having similar behavior at high temperatures: a paramagnetic regime where thermal fluctuations govern the system. At around the first crossover, the behavior of the isotropic and the anisotropic systems still matches. Here, the system crosses over from a paramagnetic regime to a regime where spins on a tetrahedron follow the ice rule. However, below this point, the behavior of the different phases is very different. In fact, we can observe either a second crossover (in the planar ice phase), indicated by a second bump in the specific heat, or a phase transition toward an ordered phase, indicated by a divergence in the specific heat. This allows us to make a distinction between the symmetry-broken phases and the non-symmetry-broken phases. For all systems, we can observe that the entropy density is lower than the Pauling estimate. Next, we analyze the five different phases in more detail and distinguish between the non-symmetry-broken phases and the symmetry-broken phases at finite temperatures.
@@ IV.2.1 Non-symmetry-broken phases
[…]
@@ IV.2.2 Symmetry-broken phases
If on any sublattice the strain-parameter $\delta_{A}/\delta_{B}<0$, a phase transition into a symmetry-broken state will occur at a finite temperature. The phase transition presents itself as a divergence in the specific heat. Due to finite-size effects, the divergence will appear as a narrow peak with a finite height in Fig. 3.
[…]
For the ferromagnetic case, where $\delta_{A}=\delta_{B}<0$, it has been shown that a transition akin to the one found in hydrogen-bonded ferroelectrics known as the KDP (potassium dihydrogen phosphate) model [63, 64] takes place, ordering the system ferromagnetically. Here, a sharp divergence of the specific heat, as well as a sharp dip in the entropy, is visible [17, 18]. We can calculate the exact critical temperature by considering the energy cost of a string excitation and the entropy of this excitation [28]. Let us consider a ground state of the system. All the spins are equally aligned on all tetrahedra, such that the global magnetization vector of all tetrahedra points in the same direction. A string excitation, now, consists of flipping a row of spins in the $z$ direction and thereby increasing the energy. Since we are introducing a partial charge in 4 tetrahedra per unit cell, for $L_{z}$ unit cells in $z$ direction, the energy for such an excitation is $\Delta U=4\delta\cdot 4L_{z}=16\delta L_{z}$. Because there are two possible ways for a string to enter each tetrahedron, and we flip $4L_{z}$ tetrahedra, the entropic gain of a string excitation is $\Delta S=\ln(2^{4L_{z}})$. So the difference in free energy for a string excitation is
@@
In summary, in all cases, the specific heat has a second signature below the first crossover at $T\approx J$. If on any sublattice there is a negative strain parameter, $\delta_{i}<0$, we expect a phase transition at a finite temperature into a symmetry-broken phase. The symmetry breaking will, however, occur only in dimensionally reduced subsystems if one of the strain parameters is non-negative. Here, the order parameter is defined on a dimensionally reduced support – a two-dimensional plane – and the symmetry that is broken is the intermediate symmetry of the plane [32, 33, 31]. This also leads to a vanishing entropy density. If both strain parameters are positive, one-dimensional lines will effectively decouple, leading to antiferromagnetic order on lines at zero temperature. Similar behavior has been observed in the anisotropic antiferromagnetic triangular lattice Ising model, where strong coupling along one lattice direction leads to effective dimensional reduction and frustrated inter-chain interactions. Despite the two-dimensional connectivity, the system may exhibit behavior akin to weakly coupled one-dimensional chains without a finite-temperature phase transition [65]. If only one strain parameter is non-zero and positive, a crossover into a regime where the two-dimensional planes decouple with an extensive ground state degeneracy occurs. Within planes, the planar ice constraint holds. This is the only phase where the entropy density stays finite even for an infinite system size.
@@
Here, we assume the systems to be at equilibrium. To reach thermal equilibrium in numerical simulations, techniques such as cluster updates or worm loop updates are employed to overcome large energy barriers. In fact, we can observe that the acceptance rate for single spin-flips is reduced significantly below the first crossover into the spin ice regime at $T\approx J$ (see App. F). For isotropic spin ice, the worm algorithm has been proven useful to overcome the long waiting times to overcome the energy barriers and equilibrate spin ice systems [15, 55, 56, 57]. The acceptance rate for worm loop updates is exactly one below the crossover into the spin ice manifold for the isotropic case. However, for the anisotropic systems, even the acceptance rate for the worm algorithm reduces close to the crossover into the respective restricted manifold. This renders equilibrating the anisotropic systems numerically very hard, as updates are correlated with extremely long waiting times that increase with increasing system size. It is, however, important to point out that these are higher-order processes that will also take extremely long in an experimental setup. It is expected that cooling the system too quickly will hide these features, and the system might then seem to behave akin to spin ice [9].
@@ V Spin Structure Factor
In a scattering event with neutron polarization parallel to $\hat{z}_{\text{sc}}$, one can define the orthonormal basis vectors for each scattering wave vector $\mathbf{q}\perp\hat{z}_{\text{sc}}$ as $\hat{x}_{\text{sc}}\equiv\hat{\mathbf{q}}$ and $\hat{y}_{\text{sc}}\equiv\hat{z}_{\text{sc}}\times\hat{x}_{\text{sc}}$. The spin vector is defined as $\hat{S}^{z}_{i,\alpha}=S^{z}_{i,\alpha}\cdot\hat{e}_{i}$, where $\alpha$ is the index referring to the four sites in the face-centered cubic (fcc) unit cell, while $i$ is the spin index, referring to the site on a single tetrahedron. The unit vectors $\hat{e}_{i}$ are the local basis vectors pointing into the center of a single tetrahedron, as defined in App. C.
@@
In polarized neutron scattering experiments, with $\hat{z}_{\text{sc}}\parallel[1\bar{1}0]$, the spin ice material Ho2Ti2O7 and Dy2Ti2O7 show particular pinch points in the $[hhl]$ plane in the SF channel, indicative of the dipolar form of the spin correlations in three dimensions, while in the NSF channel no particular features can be seen [2, 66]. Now, we want to understand how this structure factor will differ in the anisotropic phases compared to isotropic spin ice. Numerically, the spin structure factor can be obtained by sampling states via a Monte Carlo procedure or by an analytical self-consistent Gaussian approximation. We will employ both techniques and show the correspondence of the data in Fig. 4.